A Level Mathematics 9709 Paper 63, Oct/Nov 2024

Question paper and mark scheme50 marks7 questions

Solve it on ExamPal

Practise

Untimed. Write on the paper and check as you go.

Start

Exam

Timed, with a clock you set, like sitting the real paper.

Start

Opens full screen.

Papers open in

Opening the question paper

Read as textThe questions and the mark scheme as plain text. Easier to read on a phone.

Question paper

9709/63 · Oct/Nov 2024

Download PDF
1

The heights of a certain species of deer are known to have standard deviation 0.35 m. A zoologist takes a random sample of 150 of these deer and finds that the mean height of the deer in the sample is 1.42 m.

a

Calculate a 96% confidence interval for the population mean height.

[3]
b

Bubay says that 96% of deer of this species are likely to have heights that are within this confidence interval. Explain briefly whether Bubay is correct.

[1]
2

The masses, in kilograms, of small and large bags of wheat have the independent distributions N(16.0,0.4) and N(51.0,0.9) respectively. Find the probability that the total mass of 3 randomly chosen small bags is greater than the mass of one randomly chosen large bag.

[5]
3

The times, T minutes, taken by a random sample of 75 students to complete a test were noted. The results were summarised by Σt=230 and Σt2=930.

a

Calculate unbiased estimates of the population mean and variance of T.

[3]
b

You should now assume that your estimates from part (a) are the true values of the population mean and variance of T. The times taken by another random sample of 75 students were noted, and the sample mean, T‾, was found. Find the value of a such that P(T‾>a)=0.234.

[3]
4

A random variable X has probability density function f defined by f(x)={ax2−18x32⩽x⩽3,0otherwise, where a is a constant.

a

Show that a=272.

[3]
b

Show that E(X)=272ln⁡32−3.

[3]
5

The lengths, in centimetres, of worms of a certain kind are normally distributed with mean μ and standard deviation 2.3. An article in a magazine states that the value of μ is 12.7. A scientist wishes to test whether this value is correct. He measures the lengths, x cm, of a random sample of 50 worms of this kind and finds that Σx=597.1. He plans to carry out a test, at the 1% significance level, of whether the true value of μ is different from 12.7.

a

State, with a reason, whether he should use a one-tailed or a two-tailed test.

[1]
b

Carry out the test.

[5]
6

The numbers of customers arriving at service desks A and B during a 10-minute period have the independent distributions Po(1.8) and Po(2.1) respectively.

a

Find the probability that during a randomly chosen 15-minute period more than 2 customers will arrive at desk A.

[2]
b

Find the probability that during a randomly chosen 5-minute period the total number of customers arriving at both desks is less than 4.

[3]
c

An inspector waits at desk B. She wants to wait long enough to be 90% certain of seeing at least one customer arrive at the desk. Find the minimum time for which she should wait, giving your answer correct to the nearest minute.

[4]
7

The number of accidents per year on a certain road has the distribution Po(λ). In the past the value of λ was 3.3. Recently, a new speed limit was imposed and the council wishes to test whether the value of λ has decreased. The council notes the total number, X, of accidents during two randomly chosen years after the speed limit was introduced and it carries out a test at the 5% significance level.

a

Calculate the probability of a Type I error.

[4]
b

Given that X=2, carry out the test.

[3]
c

The council decides to carry out another similar test at the 5% significance level using the same hypotheses and two different randomly chosen years. Given that the true value of λ is 0.6, calculate the probability of a Type II error.

[3]
d

Using λ=0.6 and a suitable approximating distribution, find the probability that there will be more than 10 accidents in 30 years.

[4]

Mark scheme

9709/63 · Oct/Nov 2024

Download PDF

1a · 3 marks

  • B1

    z = 2.054 or 2.055

  • M1

    1.42 ± z 0.35 / √150

  • A1

    1.36 to 1.48 [m] (3 sf)

1b · 1 mark

  • B1

    No. CI is about mean, not individual values.

2 · 5 marks

  • B1

    E(S₁ + S₂ + S₃ – L) = 16 x 3 – 51 [= −3 ]

  • M1

    Var(S₁ + S₂ + S₃ – L) = 3 × 0.4 + 0.9 [= 2.1]

  • M1

    (0 − (−3)) / √'2.1' [= 2.070]

  • M1

    1 − Φ(‘2.070’)

  • A1

    = 0.0192 (3 sf)

3a · 3 marks

  • B1

    t̄ = 230 / 75 [= 3.0666… or 3.07 (3 sf)] [ 0r 46/15 ]

  • M1

    s² = 75/74 (930/75 − (230/75)²) or 1/74(930 – 230²/75)

  • A1

    = 3.0360… or 3.04 (3 sf) or = 337/111

3b · 3 marks

  • B1

    [ Φ⁻¹(1 − 0.234) ] = 0.726

  • M1

    ± (a − '3.0667') / √('3.04'/75) = ± ‘0.726’

  • A1

    a = 3.21 (3 sf)

4a · 3 marks

  • M1

    ∫₂³ (a/x² − 18/x³) dx = 1

  • A1

    [−a/x + 9/x²]₂³ = 1

  • A1

    [−a/3 + 1 + a/2 − 9/4] = 1 ⇒ a = 27/2 (AG)

4b · 3 marks

  • M1

    ∫₂³ (27/(2x) − 18/x²) dx

  • A1

    [27/2 ln x + 18/x]₂³ or [27/2 ln 2x + 18/x]₂³

  • A1

    = 27/2 ln 3 + 6 − 27/2 ln 2 − 9 = 27/2 ln(3/2) − 3 AG

5a · 1 mark

  • B1

    Two-tailed because looking for difference

5b · 5 marks

  • B1

    H₀: μ = 12.7 H₁: μ ≠ 12.7

  • M1

    (597.1/50 − 12.7) / (2.3 / √50)

  • A1

    = −2.330 Accept – 2.336 or – 2.337

  • M1

    ‘−2.330’ > −2.576 or ‘2.330’ < 2.576 or ‘0.00989’ > 0.005 or 0.0097 > 0.005

  • A1

    [Not reject H₀] There is insufficient evidence to suggest that µ is not 12.7

6a · 2 marks

  • M1

    [λ = 2.7] 1 − e⁻²·⁷(1 + 2.7 + 2.7²/2) or 1 − e⁻²·⁷(1 + 2.7 + 3.645) or 1- ( 0.06721 + 0.1815 + 0.2450 )

  • A1

    = 0.506 (3 sf)

6b · 3 marks

  • B1

    λ = 1.95

  • M1

    e⁻¹·⁹⁵(1 + 1.95 + 1.95²/2 + 1.95³/3!) or e⁻¹·⁹⁵(1 + 1.95 + 1.90125 + 1.2358) or 0.1423 + 0.2774+ 0.2705 + 0.1758

  • A1

    = 0.866

6c · 4 marks

  • M1

    1 – e⁻²·¹ˣ ⩾ 0.90 or 1 – e⁻ᵞ ⩾ 0.90

  • M1

    [e⁻²·¹ˣ ≤ 0.1] or e⁻ᵞ ≤ 0.1 −2.1x ≤ ln0.1 or −λ ≤ ln0.1 [ λ ≥ 2.3026, 2.3026/2.1 ]

  • A1

    1.096 or 10.96 accept 1.097 or 10.97

  • A1

    She must wait for at least 11 minutes

7a · 4 marks

  • B1

    λ = 6.6

  • M1

    P(X ≤ 2) = e⁻⁶·⁶(1 + 6.6 + 6.6²/2) [= 0.0400] [ < 0.05 ] or e⁻⁶·⁶(1 + 6.6 + 21.78 ) or 0.001360 + 0.008978 + 0.02963

  • B1

    P(X ≤ 3) = e⁻⁶·⁶(1 + 6.6 + 6.6²/2 + 6.6³/3!) or 0.0400 + e⁻⁶·⁶ × 6.6³/3! = 0.105 [ > 0.05 ]

  • A1

    P(Type I error) = 0.0400 (3 sf)

7b · 3 marks

  • B1

    H₀: λ = 6.6, H₁: λ < 6.6

  • M1

    [P(X ≤ 2) = 0.0400] 0.04 < 0.05

  • A1

    [Reject H₀] There is evidence to suggest that mean number of accidents has decreased

7c · 3 marks

  • M1

    P(X > 2) attempted, with any λ

  • M1

    P(X > 2) = 1 − e⁻¹·²(1 + 1.2 + 1.2²/2) or = 1 − e⁻¹·²(1 + 1.2 +0.72) or = 1 – ( 0.3012 + 0.3614 + 0.2169 )

  • A1

    0.121 (3 sf) or 0.120

7d · 4 marks

  • B1

    N(18, 18) seen or implied

  • M1

    (10.5 − 18) / √18 [= −1.768]

  • M1

    P(X > ‘−1.768’) = Φ(‘1.768’)

  • A1

    = 0.961 or 0.962 (3 sf)

Questions

  1. 1
    Question 14 marks
    Mark scheme ›
  2. 2
    Question 25 marks
    Mark scheme ›
  3. 3
    Question 36 marks
    Mark scheme ›
  4. 4
    Question 46 marks
    Mark scheme ›
  5. 5
    Question 56 marks
    Mark scheme ›
  6. 6
    Question 69 marks
    Mark scheme ›
  7. 7
    Question 714 marks
    Mark scheme ›

More 9709 papers

Other papers from Oct/Nov 2024

Solve it on ExamPalStart